OpenStudy (anonymous):

A truck is traveling 70.0 kph away from you. The driver is blowing the horn, which has a frequency of 400 Hz. The air temperature is 25°C. What is the observed frequency of the horn?

OpenStudy (aaronq):

speed of truck =70 km/h*\(\sf \dfrac{1~hr}{60~min}*\dfrac{1~min}{60~sec}*\dfrac{1000~m}{1~km}\)= 19.44 m/s speed of sound in air =343 m/s doppler formula for receding object stationary observer \(\sf f_{obs}=\dfrac{f_o*v}{v_s+v}=\dfrac{400~Hz*343~m/s}{19.44m/s+343~m/s}=378.54=379~Hz\)